# Why Knowledge Distillation Works in Generative Models: A Minimal Working Explanation

### v1

@@ Abstract
Knowledge distillation (KD) is a core component in the training and deployment of modern generative models, particularly large language models (LLMs). While its empirical benefits are well documented—enabling smaller student models to emulate the performance of much larger teachers—the underlying mechanisms by which KD improves generative quality remain poorly understood. In this work, we present a minimal working explanation of KD in generative modeling. Using a controlled simulation with mixtures of Gaussians, we demonstrate that distillation induces a trade-off between precision and recall in the student model. As the teacher distribution becomes more selective, the student concentrates more probability mass on high-likelihood regions at the expense of coverage—a behavior modulated by a single entropy-controlling parameter. We then validate this effect in a large-scale language modeling setup using the SmolLM2 family of models. Empirical results reveal the same precision–recall dynamics observed in simulation, where precision corresponds to sample quality and recall to distributional coverage. This precision–recall trade-off proves especially beneficial in scenarios where sample quality outweighs diversity, such as instruction tuning or downstream generation. Our analysis provides a simple and general explanation for the effectiveness of KD in generative modeling.
@@ 1 Introduction
Knowledge distillation (KD) has become a foundational technique in modern machine learning. Originally introduced as a method to compress large classification models by transferring knowledge from a teacher to a smaller student model (hinton2015distilling, ), KD has since proven effective in improving generalization, stabilizing training, and aligning model behavior, in several domains (romero2014fitnets, ; li2017learning, ; tian2019contrastive, ). Its influence is especially prominent in generative modeling, where KD is now a standard component in the training and deployment of large language models (LLMs). From neural machine translation (kim2016sequence, ) to the latest instruction-tuned LLMs (ko2024distillm, ; abouelenin2025phi, ; grattafiori2024llama, ), distillation enables smaller models to generate coherent and high-quality text by mimicking the output distributions of larger models. Despite its ubiquity, however, the mechanisms by which KD improves generative performance remain poorly understood, particularly in how distillation shapes the generative behavior of the student model.
@@
While several studies have attempted to explain KD in the context of classification—emphasizing representation alignment, label smoothing effects, or decision boundary refinement (phuong2019towards, ; muller2019does, ; stanton2021does, )—these analyses do not generalize naturally to autoregressive generative models. In particular, there is little theoretical understanding of how KD enables smaller generative models to achieve performance comparable to much larger models, even with significantly reduced capacity. Why do student models trained via KD generate higher-quality outputs than those trained via maximum likelihood? What inductive bias does the teacher introduce during distillation that improves sample quality? These questions remain open, despite KD’s critical role in the development of state-of-the-art LLMs.
@@
In this paper, we provide a minimal working explanation of knowledge distillation in generative models by analyzing how distillation reshapes the student’s learned distribution. We begin with a controlled simulation using mixtures of Gaussians and show that distillation induces a trade-off between precision and recall: as the teacher distribution becomes lower in entropy, the student concentrates more probability mass on high-likelihood regions while sacrificing coverage. This behavior is governed by a single temperature-like parameter that controls the selectivity of the teacher’s output. We then validate this insight in a large-scale language modeling setting. Specifically, we use SmolLM2 1.7B (allal2025smollm2, ) as the ground-truth distribution to generate samples, on which we pretrain a 360M teacher model. This teacher is subsequently distilled into a 135M student model. Empirical results corroborate our theoretical predictions: as the teacher becomes more selective, the student produces sharper generations (i.e., higher precision) at the cost of reduced recall—mirroring the trade-off observed in the Gaussian mixture setup. These findings suggest that distillation enables smaller generative models to concentrate probability mass on high-density regions of the output space, effectively producing sharper and more fluent generations. This trade-off is especially desirable in practical scenarios such as instruction tuning or task-specific generation, where sample quality is prioritized over full coverage. Our contributions are summarized as follows:
@@
  • We provide a minimal working explanation of knowledge distillation in generative models, highlighting a precision–recall trade-off that emerges from teacher entropy.
@@
  • We validate this mechanism in large-scale language models, where precision–recall dynamics observed in simulation are replicated via multistage distillation from SmoLLM2 1.7B to 360M to 135M.
@@ 2 Related Work
Knowledge distillation.   Knowledge distillation (KD) was originally introduced as a model compression technique that transfers knowledge from a cumbersome teacher model into a smaller, faster student hinton2015distilling . By matching the teacher’s softened output probabilities, the student can capture richer inter-class similarity patterns, leading to improved generalization and efficiency ba2014deep ; urban2016deep . This foundational idea has inspired a range of extensions. Born-Again Networks demonstrate that students can outperform their teachers through repeated distillation furlanello2018born , while FitNets leverage intermediate feature representations to guide deeper students during training romero2014fitnets . Recent work further suggests that KD can improve generalization even when the teacher and student have identical capacities huang2021revisiting ; stanton2021does , indicating that the utility of KD extends well beyond model compression.
@@
Broader use of KD across domains.   KD has become a central component in modern NLP systems and generative modeling. In neural machine translation, sequence-level KD was proposed to improve generation quality while reducing model size kim2016sequence . More recently, state-of-the-art LLM frameworks such as DISTILLM ko2024distillm , Phi-4-Mini abouelenin2025phi , and LLaMA 3 grattafiori2024llama have adopted KD as a core technique for post-training alignment and efficient deployment. Beyond language models, KD has also been applied in diverse domains such as self-supervised learning tian2019contrastive , speech recognition chebotar2016distilling , and continual learning li2017learning ; cha2024regularizing , often to transfer behaviors from larger or prior models into smaller or adaptive ones. Despite its broad adoption, most implementations of KD treat it as a black-box heuristic—lacking a clear understanding of its internal mechanisms, particularly in generative settings.
@@
Why does KD work?   A number of theoretical efforts have attempted to explain why KD works. In linear models, soft targets have been shown to produce more robust and smoother decision boundaries phuong2019towards . KD has also been connected to label smoothing, suggesting that it introduces an implicit regularization effect muller2019does . Empirical studies have confirmed that KD improves generalization across a variety of conditions, even when teacher and student architectures are identical stanton2021does . Additional work frames KD as a trade-off between knowledge inheritance and exploration huang2021revisiting . However, these interpretations have been largely developed in the context of classification tasks. There remains a significant gap in understanding how KD influences the generative behavior of models, particularly autoregressive language models—a gap that our work aims to address.
While prior work has provided valuable insights into the mechanisms of knowledge distillation, most analyses have focused on classification tasks or relied on task-specific assumptions. In contrast, our work investigates KD in the context of generative modeling and seeks to understand how the teacher’s selectivity influences the student’s learned distribution.
3 Distillation in a Generative Model: with a Mixture of Gaussians
@@
To understand the core effect of knowledge distillation in a generative modeling, we introduce a simple yet expressive setup based on mixtures of Gaussians. This construction allows us to precisely control the complexity of the data distribution and quantify how the student model responds to different teacher behaviors. By modulating the entropy of the teacher distribution through a temperature-like parameter $\beta$, we adjust how selectively the teacher emphasizes certain regions of the data space. This minimal setting reveals a key trade-off between precision and recall in the student model, providing insight into how distillation alters the student’s learned distribution.
3.1 Distillation as a way to balance precision and recall
@@
In the extreme case of $K^{\prime}=K$, $\sigma$ would ideally be a permutation operator, and all the parameters are well recovered up to some noise due to the finite number of samples within $D$. This noise is an important aspect here, and pretty much tells us that even if $\alpha_{k}=\frac{1}{K}$, we would never end up with the uniform $\alpha^{\prime}_{k}$. In the other extreme case of $K^{\prime}=1$, $\sigma(k^{\prime})=1$ for all $k^{\prime}$, and we end up with
@@
$\geq\sum_{k^{\prime}=1}^{K^{\prime}}\sum_{k^{\prime\prime}=1}^{K^{\prime\prime}}\underbrace{\alpha_{k^{\prime}}^{\prime}(\beta)\alpha_{k^{\prime\prime}}^{\prime\prime}}_{\text{(a')}}\underbrace{\int\mathcal{N}(x;\mu_{k^{\prime}}^{\prime},\Sigma^{\prime}_{k^{\prime}})\log\mathcal{N}(x;\mu_{k^{\prime\prime}}^{\prime\prime},\Sigma_{k^{\prime\prime}}^{\prime\prime})\mathrm{d}x}_{\text{(b')}},$
@@
This formulation reveals how the entropy of the teacher—controlled by $\beta$—governs the selectivity of the student’s focus. Larger $\beta$ values lead to a more peaked teacher distribution, which in turn biases the student toward modeling fewer but sharper modes.
@@
This quantity closely correlates with the number of emphasized teacher components, i.e., $\sum_{k^{\prime}=1}^{K^{\prime}}\mathds{1}(\alpha^{\prime}_{k^{\prime}}\geq 1-\epsilon)$. Crucially, this difficulty is directly modulated by the temperature parameter $\beta$, which controls the entropy of the teacher distribution. Higher $\beta$ values yield more selective (lower entropy) teachers, concentrating training signals on a smaller subset of $p^{*}$ and thereby reducing the student’s modeling burden.
@@
$\mathrm{Precision}(\beta)=\mathbb{E}_{p^{\prime\prime}(x;\theta^{\prime\prime},\beta)}\left[\log p^{*}(x;\theta^{*})\right].$ (3)
@@
Intuitively, precision measures how well the student’s samples are supported under the true distribution $p^{*}$, while recall quantifies how thoroughly the student covers the modes of $p^{*}$. As $\beta\to\infty$, the teacher distribution becomes increasingly selective, concentrating mass on a small subset of high-density regions. The student, in turn, learns to model these regions accurately—leading to high precision but reduced recall. Conversely, when $\beta\approx 1$, the teacher approximates the full support of $p^{*}$, and the student is encouraged to match the entire distribution. This maximizes recall but often comes at the cost of lower precision due to limited capacity.
By modulating $\beta$, distillation provides a simple mechanism to control this trade-off. Our analysis reveals that this precision–recall dynamic emerges naturally from the structure of the distillation objective and offers a principled explanation for why distillation can improve sample quality in generative models, even under constrained capacity.
Figure 1: (a) The contours correspond to the probability density of, and the dots are the samples drawn from $p^{*}$. (b) The contours correspond to the student model trained directly without distillation.
Figure 2: (a) Contour plot of the teacher distribution $p^{\prime}$, with samples from the true distribution $p^{*}$ overlaid as blue dots. (b) Samples (green) drawn from the $\beta$-modulated teacher distribution $p^{\prime}(x;\theta^{\prime},\beta=100)$, showing strong concentration on the bottom three modes of $p^{*}$. (c) Contours of the student models: the dashed black contour corresponds to a student trained directly on $p^{*}$ samples, while the green contour represents a student trained on teacher samples (distillation). The distilled student clearly focuses on a narrower region emphasized by the teacher. This setup corresponds to a low-difficulty case: with $\beta=100$, the teacher concentrates on a single dominant component ($\alpha^{\prime}_{k}\approx 1$), and the student has just enough capacity ($K^{\prime\prime}=1$) to match it, yielding a difficulty of 0.
@@ 3.2 Simulation
To empirically illustrate the mechanisms discussed above, we construct a toy generative task using a mixture of Gaussians. The ground-truth distribution $p^{*}$ consists of eight well-separated components arranged in a rectangular grid (Fig. 1a). We begin by fitting a student model composed of a single Gaussian directly on samples from $p^{*}$. As shown in Fig. 1b, this model places its density around the center of the space—an area with near-zero mass under $p^{*}$—demonstrating the difficulty of approximating a complex multimodal distribution with a simple model under standard MLE training.
@@
We then sample training data from a temperature-modulated version of the teacher, $p^{\prime}(x;\beta=100)$. At this high value of $\beta$, sampling concentrates on the teacher component with the largest weight, yielding data that primarily covers the bottom three modes of $p^{*}$ (Fig. 2a–b).
A new student model is then trained on these samples. Fig. 2c compares the density contours of the directly trained student (dashed black) and the distilled student (green). The distilled model clearly focuses on a smaller region of $p^{*}$ where the teacher emphasized high-density modes, while ignoring other parts of the support.
@@
  • Direct student: Precision = $-20.26$, Recall = $-2.64$
  • Distilled student ($\beta=100$): Precision = $-0.70$, Recall = $-42.45$
These results confirm our theoretical prediction: as the teacher becomes more selective (higher $\beta$), the student learns to place greater mass on high-likelihood regions at the cost of covering the full distribution. In conclusion, knowledge distillation thus enables simpler models to generate sharper, more concentrated outputs—a desirable property when sample quality is prioritized over full coverage.

### v3

@@ Abstract
Knowledge distillation (KD) is a core component in the training and deployment of modern generative models, particularly large language models (LLMs). While its empirical benefits are well documented—enabling smaller student models to emulate the performance of much larger teachers—the underlying mechanisms by which KD improves generative quality remain poorly understood. In this work, we present a minimal working explanation of KD in generative modeling. Using a controlled simulation with mixtures of Gaussians, we demonstrate that distillation induces a trade-off between precision and recall in the student model. As the teacher distribution becomes more selective, the student concentrates more probability mass on high-likelihood regions at the expense of coverage, which is a behavior modulated by a single entropy-controlling parameter. We then validate this effect in a large-scale language modeling setup using the SmolLM2 family of models. Empirical results reveal the same precision-recall dynamics observed in simulation, where precision corresponds to sample quality and recall to distributional coverage. This precision-recall trade-off in LLMs is found to be especially beneficial in scenarios where sample quality is more important than diversity, such as instruction tuning or downstream generation. Our analysis provides a simple and general explanation for the effectiveness of KD in generative modeling.
@@ 1 Introduction
Knowledge distillation (KD) has become a foundational technique in modern machine learning. Originally introduced as a method to compress large classification models by transferring knowledge from a teacher to a smaller student model (Hinton et al., 2015), KD has since proven effective in improving generalization and aligning model behavior in several domains (Romero et al., 2014; Li and Hoiem, 2017; Tian et al., 2019). Its influence is especially prominent in generative modeling, where KD is now a standard component in the training and deployment of large language models (LLMs). From neural machine translation (Kim and Rush, 2016) to the latest instruction-tuned LLMs (Ko et al., 2024; Abouelenin et al., 2025; Grattafiori et al., 2024), distillation enables smaller models to generate coherent and high-quality text by mimicking the output distributions of larger models Gu et al. (2024); Agarwal et al. (2024); Wen et al. (2023). Despite its ubiquity, however, the mechanisms by which KD improves generative performance remain poorly understood, particularly in how distillation shapes the generative behavior of the student model.
@@
While several studies have attempted to explain KD in the context of classification, emphasizing representation alignment, label smoothing effects, or decision boundary refinement (Phuong and Lampert, 2019; Müller et al., 2019; Stanton et al., 2021), these analyses do not generalize naturally to autoregressive generative models. In particular, there is little theoretical understanding of how KD enables smaller generative models to achieve performance comparable to much larger models, even with significantly reduced capacity. Why do student models trained via KD generate higher-quality outputs than their maximum likelihood-trained counterparts? What inductive bias does the teacher introduce during distillation that improves sample quality? These questions remain open, despite KD’s critical role in the development of high-performing yet efficient LLMs.
@@
In this paper, we provide a minimal working explanation of knowledge distillation in generative models by analyzing how distillation reshapes the student’s learned distribution. We begin with a controlled simulation using Gaussian mixtures and show that distillation induces a trade-off between precision and recall: as the teacher distribution becomes lower in entropy, the student concentrates more probability mass on high-likelihood regions while sacrificing coverage. This behavior is governed by a single temperature-like parameter that controls the selectivity of the teacher’s output. We then validate this insight in a large-scale language modeling setting. Specifically, we use SmolLM2 1.7B (Allal et al., 2025) as the ground-truth distribution to generate samples, on which we pretrain a 360M teacher model. This teacher is subsequently distilled into a 135M student model. Empirical results corroborate our theoretical predictions: as the teacher becomes more selective, the student produces sharper generations (i.e., higher precision) at the cost of reduced recall—mirroring the trade-off observed in the Gaussian mixture framework. These findings suggest that distillation enables smaller generative models to concentrate probability mass on high-density regions of the output space, effectively producing sharper and more fluent generations. This trade-off is especially desirable in practical scenarios such as instruction tuning or task-specific generation, where sample quality is prioritized over full coverage. Our contributions are summarized as follows:
@@
  • We provide a minimal working explanation of knowledge distillation in generative models, highlighting a precision-recall trade-off that emerges from teacher entropy.
@@
  • We validate this mechanism in large-scale language models, where precision-recall dynamics observed in simulation are replicated via multistage distillation from SmolLM2 1.7B to 360M to 135M.
@@ 2 Related Work
Knowledge distillation. Knowledge distillation (KD) was originally introduced as a model compression technique that transfers knowledge from a large teacher model into a smaller student Hinton et al. (2015). By matching the teacher’s softened output probabilities, the student can capture richer inter-class similarity patterns, leading to improved generalization and efficiency Ba and Caruana (2014); Urban et al. (2016). This foundational idea has inspired a range of extensions. Born-Again Networks Furlanello et al. (2018) demonstrate that students can outperform their teachers through repeated distillation, while FitNets Romero et al. (2014) leverage intermediate feature representations to guide deeper students during training. Recent work further suggests that KD can improve generalization even when the teacher and student have identical capacities Huang et al. (2021); Stanton et al. (2021), indicating that the utility of KD extends well beyond model compression.
@@
Broader use of KD across domains. KD has become a central component in modern NLP systems and generative modeling. In neural machine translation, sequence-level KD was proposed to improve generation quality while reducing model size Kim and Rush (2016). More recently, state-of-the-art LLM frameworks such as DISTILLM Ko et al. (2024), Phi-4-Mini Abouelenin et al. (2025), and LLaMA 3 Grattafiori et al. (2024) have adopted KD as a core technique for post-training alignment and efficient deployment. Beyond language models, KD has also been applied in diverse domains such as self-supervised learning Tian et al. (2019), speech recognition Chebotar and Waters (2016), and continual learning Li and Hoiem (2017); Cha et al. (2024), often to transfer behaviors from larger or prior models into smaller or adaptive ones. Despite its broad adoption, most implementations of KD treat it as a black-box heuristic, lacking a clear understanding of its internal mechanisms, particularly in generative settings.
@@
Why does KD work? A number of theoretical efforts have attempted to explain why KD works. In linear models, soft targets have been shown to produce more robust and smoother decision boundaries Phuong and Lampert (2019). KD has also been connected to label smoothing, suggesting that it introduces an implicit regularization effect Müller et al. (2019). Empirical studies have confirmed that KD improves generalization across a variety of conditions, even when teacher and student architectures are identical Stanton et al. (2021). Additional work frames KD as a trade-off between knowledge inheritance and exploration Huang et al. (2021). However, these interpretations have been largely developed in the context of classification tasks. There remains a significant gap in understanding how KD influences the generative behavior of models, particularly autoregressive language models—a gap that our work aims to address.
Recent studies on KD for generative models have explored related concepts like mode-seeking behavior Gu et al. (2024); Agarwal et al. (2024); Wen et al. (2023). Often, achieving such behavior involved proposing specialized objectives, like reverse KL divergence Gu et al. (2024); Wen et al. (2023) or tailored loss functions Agarwal et al. (2024). In contrast, our work provides a different perspective by demonstrating how this precision-enhancing effect naturally arises within the standard forward KL divergence framework, simply by controlling the teacher distribution’s selectivity. Furthermore, we analyze this phenomenon not just between teacher and student, but within a broader three-stage framework (ground truth $\rightarrow$ teacher $\rightarrow$ student), offering a fundamental, distribution-level explanation based on a precision-recall trade-off for why common KD practices improve sample quality relative to the original data distribution.
3 Distillation in Generative Modeling: Analysis via Gaussian Mixtures
@@
To understand the core effect of knowledge distillation in generative modeling, we introduce a simple yet expressive setup based on mixtures of Gaussians. This construction allows us to precisely control the complexity of the data distribution and quantify how the student model responds to different teacher behaviors. By modulating the entropy of the teacher distribution through a temperature-like parameter $\beta$, we adjust how selectively the teacher emphasizes certain regions of the data space. This minimal setting reveals a key trade-off between precision and recall in the student model, providing insight into how distillation alters the student’s learned distribution.
3.1 The precision-recall trade-off induced by distillation
@@
In the extreme case of $K^{\prime}=K$, $\sigma$ would ideally be a permutation operator, and all the parameters are well recovered up to some noise due to the finite number of samples within $D$. This noise is an important aspect here, which implies that even if $\alpha_{k}=\frac{1}{K}$, we would never end up with the uniform $\alpha^{\prime}_{k}$. In the other extreme case of $K^{\prime}=1$, $\sigma(k^{\prime})=1$ for all $k^{\prime}$, and we end up with
@@
$\geq\sum_{k^{\prime}=1}^{K^{\prime}}\sum_{k^{\prime\prime}=1}^{K^{\prime\prime}}\underbrace{\alpha_{k^{\prime}}^{\prime}(\beta)\alpha_{k^{\prime\prime}}^{\prime\prime}}_{\text{(a')}}\underbrace{\int\mathcal{N}(x;\mu_{k^{\prime}}^{\prime},\Sigma^{\prime}_{k^{\prime}})\log\mathcal{N}(x;\mu_{k^{\prime\prime}}^{\prime\prime},\Sigma_{k^{\prime\prime}}^{\prime\prime})\mathrm{d}x}_{\text{(b')}}.$
@@
This formulation reveals how the entropy of the teacher, controlled by $\beta$, governs the selectivity of the student’s focus. Larger $\beta$ values lead to a more peaked teacher distribution, which in turn biases the student toward modeling fewer but sharper modes.
@@
This quantity closely correlates with the number of emphasized teacher components, i.e., $\sum_{k^{\prime}=1}^{K^{\prime}}\mathds{1}(\alpha^{\prime}_{k^{\prime}}\geq 1-\epsilon)$. Crucially, this difficulty is directly modulated by the temperature-like parameter $\beta$, which controls the entropy of the teacher distribution. Higher $\beta$ values yield more selective (lower entropy) teachers, concentrating training signals on a smaller subset of $p^{*}$ and thereby reducing the student’s modeling burden.
Figure 1: (a) Contours show the probability density of the ground-truth distribution $p^{*}$, with dots representing samples drawn from it. (b) The contours correspond to the student model trained directly without distillation.
@@
$\mathrm{Precision}(\beta)=\mathbb{E}_{p^{\prime\prime}(x;\theta^{\prime\prime},\beta)}\left[\log p^{*}(x;\theta^{*})\right],$ (3)
@@
Intuitively, precision measures how well samples generated by the student are supported under the true distribution $p^{*}$, while recall quantifies how thoroughly the student covers the modes of $p^{*}$. As $\beta\to\infty$, the teacher distribution becomes increasingly selective, concentrating mass on a small subset of high-density regions. The student, in turn, learns to model these regions accurately—leading to high precision but reduced recall. Conversely, when $\beta\approx 1$, the teacher approximates the full support of $p^{*}$, and the student is encouraged to match the entire distribution. This maximizes recall but often comes at the cost of lower precision due to limited capacity.
By modulating $\beta$, distillation provides a simple mechanism to control this trade-off. Our analysis reveals that this precision-recall dynamic emerges naturally from the structure of the distillation objective and offers a principled explanation for why distillation can improve sample quality in generative models, even under constrained capacity.
Figure 2: (a) Contour plot of the teacher distribution $p^{\prime}$, with samples from the true distribution $p^{*}$ overlaid as blue dots. (b) Samples (green) drawn from the $\beta$-modulated teacher distribution $p^{\prime}(x;\theta^{\prime},\beta=100)$, showing strong concentration on the bottom three modes of $p^{*}$. (c) Contours of the student models: the dashed black contour corresponds to a student trained directly on $p^{*}$ samples, while the green contour represents a student trained on teacher samples (distillation). The distilled student clearly focuses on a narrower region emphasized by the teacher. This setup corresponds to a low-difficulty case: with $\beta=100$, the teacher concentrates on a single dominant component ($\alpha^{\prime}_{k}\approx 1$), and the student has just enough capacity ($K^{\prime\prime}=1$) to match it, resulting in the Difficulty measure close to 0.
@@ 3.2 Simulation
To empirically illustrate the mechanisms discussed above, we construct a toy generative task using a mixture of Gaussians. The ground-truth distribution $p^{*}$ consists of eight well-separated components arranged in a rectangular grid (Fig. 1(a)). We begin by fitting a student model composed of a single Gaussian directly on samples from $p^{*}$. As shown in Fig. 1(b), this model places its density around the center of the space, an area with near-zero mass under $p^{*}$, demonstrating the difficulty of approximating a complex multimodal distribution with a simple model under standard MLE training.
@@
We then sample training data from a temperature-modulated version of the teacher, $p^{\prime}(x;\beta=100)$. At this high value of $\beta$, sampling concentrates on the teacher component with the largest weight, yielding data that primarily covers the bottom three modes of $p^{*}$ (Fig. 2(a) and (b)).
A new student model is then trained on these samples. Fig. 2(c) compares the density contours of the directly trained student (dashed black) and the distilled student (green). The distilled model clearly focuses on a smaller region of $p^{*}$ where the teacher emphasized high-density modes, while ignoring other parts of the support.
@@
  • Direct student: Precision = 20.26-20.26, Recall = 2.64-2.64
  • Distilled student ($\beta=100$): Precision = 0.70-0.70, Recall = 42.45-42.45
These results confirm our theoretical prediction: as the teacher becomes more selective (higher $\beta$), the student learns to place greater mass on high-likelihood regions at the cost of covering the full distribution. In conclusion, knowledge distillation enables simpler models to generate sharper, more concentrated outputs—a desirable property when sample quality is prioritized over full coverage.